# Lower Bound for The Number of Zeros in The Character Table of The Symmetric Group

### v3

@@ 3. Preliminaries
In [3], Erdös and Lehner proved that
@@
Additionally, they showed that
$1-S_{1}+S_{2}-\cdots-S_{2k-1}\leq\frac{p_{t}(N)}{p(N)}\leq 1-S_{1}+S_{2}-\cdots+S_{2k},$
where
@@
Using Rademacher’s formula [14] for the first sum, we obtain
@@
Since $t=C^{-1}\sqrt{N}\log N+x\sqrt{N}$, we approximate $\sqrt{N-t-r}=\sqrt{N}-\frac{1}{2\sqrt{N}}(t+r)+O\left(\frac{t^{2}}{N^{\frac{3}{2}}}\right)$. As $x\ll N^{\frac{1}{4}-\epsilon}$, we have $\exp\left(O\left(\frac{t^{2}}{N^{\frac{3}{2}}}\right)\right)=1+O\left(\frac{t^{2}}{N^{\frac{3}{2}}}\right)$. Now we simplify $I_{1}$ as follows
@@
The second sum tends to zero as NN becomes large. Therefore,
@@
We now proceed to prove our main theorem using the inequality $Z(N)\geq\sum_{t=1}^{N}c_{t}(N)p_{t}(N-t)$ from (1.1).
@@
Here we choose
$T_{1}$ $=\frac{\sqrt{6}}{2\pi}\sqrt{N}(\log N)\left(1+\frac{1}{2B}\right)\text{and}$
$T_{2}$ $=\frac{\sqrt{6}}{2\pi}\sqrt{N}(\log N)\left(1+\frac{1}{B}\right),$
where $B$ is defined by the relation $N^{\frac{1}{2B}}=\frac{\sqrt{6}}{2\pi}\log N$.
By Corollary 3.7, it follows that for $t\geq T_{1}$ we have
$c_{t}(N)\geq\;$ $p(N)\exp\left(-t\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)\right)\left(1+O\left(N^{-\frac{1}{2}}\log N\right)\right)$
(4.2) $=\;$ $p(N)\left(\sum_{j=0}^{\infty}(-1)^{j}\frac{t^{j}}{j!}\exp\left(-\frac{\pi tj}{\sqrt{6N}}\right)\right)\left(1+O\left(N^{-\frac{1}{2}}\log N\right)\right).$
We will first obtain a lower bound for $A_{2}$ and then for $A_{1}$. Note, by using Lemma 3.8 and Rademacher’s formula[14], we get
$p_{t}(N-t)=$ $p(N-t)\left(1+O\left(\frac{1}{\log N}\right)\right)$
(4.3) $=$ $p(N)\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)\left(1+O\left(\frac{1}{\log N}\right)\right),$
for all $t\geq T_{2}$. We plug in $c_{t}(N)$ from (4.2) and the above value of $p_{t}(N-t)$ in the expression for $A_{2}$, and get
$A_{2}\geq\;$ $p(N)\left(\sum_{t=T_{2}+1}^{N}p_{t}(N-t)+p(N)\sum_{j=1}^{\infty}\sum_{t=T_{2}+1}^{N}(-1)^{j}\frac{t^{j}}{j!}\exp\left(-\frac{\pi t(j+1)}{\sqrt{6N}}\right)\right)$
$\hskip 233.3125pt\times\left(1+O\left(\frac{1}{\log N}\right)\right)$
$=\;$ $p(N)\left(\big{(}p(N)-p_{T_{2}}(N)\big{)}+p(N)\sum_{j=1}^{\infty}\sum_{t=T_{2}+1}^{N}(-1)^{j}\frac{t^{j}}{j!}\exp\left(-\frac{\pi t(j+1)}{\sqrt{6N}}\right)\right)$
(4.4) $\hskip 233.3125pt\times\left(1+O\left(\frac{1}{\log N}\right)\right).$
Again, we use Lemma 3.8 at $t=T_{2}$ to obtain
(4.5) $p(N)-p_{T_{2}}(N)=\frac{2p(N)}{\log N}\left(1+O\left(\frac{1}{\log N}\right)\right).$
Further, $\sum_{t=T_{2}+1}^{N}$ in (4.4) can be simplified using the Abel summation
$\sum_{T_{2}<t\leq N}t^{j}\exp\left(-\frac{\pi t(j+1)}{\sqrt{6N}}\right)=$ $N^{j}\frac{\exp\left(-\frac{\pi(j+1)}{\sqrt{6N}}\right)-\exp\left(-\frac{\pi(N+1)(j+1)}{\sqrt{6N}}\right)}{1-\exp\left(-\frac{\pi(j+1)}{\sqrt{6N}}\right)}$
@@
- $j\int_{T_{2}}^{N}u^{j-1}\frac{\exp\left(-\frac{\pi(j+1)}{\sqrt{6N}}\right)-\exp\left(-\frac{\pi(u+1)(j+1)}{\sqrt{6N}}\right)}{1-\exp\left(-\frac{\pi(j+1)}{\sqrt{6N}}\right)}du$
(4.6) $=$ $\frac{2}{(j+1)\log N}\left(1+O\left(\frac{\log\log N}{\log N}\right)\right)+E_{j},$
where $|E_{j}|\leq\frac{L}{(j+1)(\log N)^{2}}$, for some positive constant $L$ independent of $N$ and $j$.
Now the estimates from (4.5) and (4.6) simplifies (4.4), and gives the following lower bound for $A_{2}$
(4.7) $A_{2}\geq$ $\left(\frac{2p(N)^{2}}{\log N}+\frac{2p(N)^{2}}{\log N}\sum_{j=1}^{\infty}\frac{(-1)^{j}}{(j+1)!}\right)\left(1+O\left(\frac{\log\log N}{\log N}\right)\right)$
$=$ $\left(2-\frac{2}{e}\right)\frac{p(N)^{2}}{\log N}\left(1+O\left(\frac{\log\log N}{\log N}\right)\right).$
Next, we calculate the lower bound for $A_{1}$. We again use (4.2) and (4.3), which are also valid for $T_{1}\leq t\leq T_{2}$, and obtain:
$\sum_{t=T_{1}+1}^{T_{2}}c_{t}(N)p_{t}(N-t)$ $=\left(\sum_{t=T_{1}+1}^{T_{2}}c_{t}(N)p(N-t)\right)\left(1+O\left(\frac{1}{\log N}\right)\right)$
(4.8) $\geq p(N)^{2}\left(\sum_{t=T_{1}+1}^{T_{2}}\exp\left(-t\exp\left(-\tfrac{\pi t}{\sqrt{6N}}\right)\right)\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)\right)$
$\hskip 159.3356pt\times\left(1+O\left(\frac{1}{\log N}\right)\right).$
@@
$\hskip 34.14322pt+G(T_{2})([T_{2}]-T_{1})-G(T_{1})([T_{1}]-T_{1})$
(4.9) $=$ $\int_{T_{1}}^{T_{2}}G(t)\,dt+O\left(N^{-\frac{1}{2}}\right).$
Changing the variable to
$u(t)=\exp\left(-\frac{\pi t}{\sqrt{6N}}\right),$
@@
Since $F^{\prime}(u)>0$ for $u\in[u(T_{2}),u(T_{1})]$, the function $F(u)$ is increasing in this interval, and thus the integral dominates near the lower endpoint $u(T_{2})$.
By the Laplace method for endpoint maxima [5], we obtain
@@
Using (4.8), (4.9) and (4.10) in (4.1), we have
(4.11) $A_{1}\geq\frac{2p(N)^{2}}{e\log N}\left(1+O\left(\frac{\log\log N}{\log N}\right)\right).$
Finally, we obtain our required lower bound for $Z(N)$ from (4.1), using the estimates for $A_{1}$ and $A_{2}$ from (4.11) and (4.7):
$Z(N)\geq\frac{2\,p(N)^{2}}{\log N}\left(1+O\left(\frac{\log\log N}{\log N}\right)\right).$
@@
By the Murnaghan-Nakayama rule 3.1 and Lemma 3.2,
(4.12) $Z_{t}(N)\geq c_{t}(N)\,p(N-t).$
Rademacher’s explicit result [14] for the partition function p(Nt)p(N-t) is given by
(4.13) $p(N-t)=\frac{1}{4(N-t)\sqrt{3}}\exp\left(\frac{2\pi}{\sqrt{6}}\sqrt{N-t}\right)\left(1+O\left((N-t)^{-1/2}\right)\right).$
Combining (4.12) and (4.13) with Proposition 3.6, we complete the proof of Theorem 1.3. ∎

### v4

@@ 3. Preliminaries
Erdös and Lehner [3] expressed $p_{t}(N)$ as follows
@@
We may observe
$1-S_{1}+S_{2}-\cdots-S_{2k-1}\leq\frac{p_{t}(N)}{p(N)}\leq 1-S_{1}+S_{2}-\cdots+S_{2k}.$
Further
@@
We simplify the above sum using the formula (1.1). It can be easily verified that $I_{2}$ tends to zero as $N\to\infty$. For $I_{1}$, note
@@
Since $t=C^{-1}\sqrt{N}\log N+x\sqrt{N}$, we approximate $\sqrt{N-t-r}=\sqrt{N}-\frac{1}{2\sqrt{N}}(t+r)+O\left(\frac{t^{2}}{N^{\frac{3}{2}}}\right)$. As $x\ll N^{\frac{1}{4}-\gamma}$, we have $\exp\left(O\left(\frac{t^{2}}{N^{\frac{3}{2}}}\right)\right)=1+O\left(\frac{t^{2}}{N^{\frac{3}{2}}}\right)$. Now we simplify $I_{1}$ as follows
@@
Therefore,
@@
We now proceed to prove our main theorem.
@@
Here, we have chosen
(4.2) $T_{1}$ $=\frac{\sqrt{6}}{2\pi}\sqrt{N}(\log N)\left(1+\frac{1}{2B}\right)\,\,\,\,\text{and }T_{2}$ $=\frac{\sqrt{6}}{2\pi}\sqrt{N}(\log N)\left(1+\frac{1}{B}\right),$
where $B$ is defined by the relation
$N^{\frac{1}{2B}}=\frac{\sqrt{6}}{2\pi}\log N.$
Before we obtain bounds for $A_{1}$ and $A_{2}$, we need the following estimates of $c_{t}(N)$ and $p_{t}(N-t)$. By part $(iv)$ of Proposition 3.7,
(4.3) $c_{t}(N)\geq\;p(N)\exp\left(-t\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)\right)\left(1+O\left(N^{-\frac{1}{2}}(\log N)^{4}\right)\right)$
(4.4) $=\;$ $p(N)\left(\sum_{j=0}^{\infty}(-1)^{j}\frac{t^{j}}{j!}\exp\left(-\frac{\pi tj}{\sqrt{6N}}\right)\right)\left(1+O\left(N^{-\frac{1}{2}}(\log N)^{4}\right)\right),\,\,\text{for }t\geq T_{1}.$
Note, by using Lemma 3.8 and equation (1.1), we get
(4.5) $p_{t}(N-t)=$ $p(N-t)\left(1+O\left(\frac{1}{\sqrt{\log N}}\right)\right)$
$=$ $p(N)\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)\left(1+O\left(\frac{1}{\sqrt{\log N}}\right)\right),$
for all $T_{1}\leq t\leq N^{\frac{3}{4}-\delta}=M$ with $0<\delta<\frac{1}{4}$. Now we will obtain the following lower bounds for $A_{2}$ and $A_{1}$.
Claim 1.
$A_{2}\geq\left(2-\frac{2}{e}\right)\frac{p(N)^{2}}{\log N}\left(1+O\left(\frac{1}{\sqrt{\log N}}\right)\right).$
Claim 2.
$A_{1}\geq\frac{2p(N)^{2}}{e\log N}\left(1+O\left(\frac{1}{\sqrt{\log N}}\right)\right).$
First, we obtain the above lower bound for $A_{2}$.
Proof of Claim 1.
We plug in $c_{t}(N)$ from (4.4) and $p_{t}(N-t)$ from (4.5) in the expression for $A_{2}$, and get
$A_{2}\geq\;$ $p(N)\left(\sum_{t=T_{2}+1}^{N}p_{t}(N-t)+\sum_{j=1}^{\infty}\sum_{t=T_{2}+1}^{N}p_{t}(N-t)(-1)^{j}\frac{t^{j}}{j!}\exp\left(-\frac{\pi tj}{\sqrt{6N}}\right)\right)$
$\hskip 204.85974pt\times\left(1+O\left(N^{-\frac{1}{2}}(\log N)^{4}\right)\right)$
$=$ $p(N)\left(\big(p(N)-p_{T_{2}}(N)\big)+p(N)\sum_{j=1}^{\infty}\sum_{t=T_{2}+1}^{M}(-1)^{j}\frac{t^{j}}{j!}\exp\left(-\frac{\pi t(j+1)}{\sqrt{6N}}\right)\right)$
$\hskip 204.85974pt\times\left(1+O\left(\frac{1}{\sqrt{\log N}}\right)\right)$
(4.6) $+$ $p(N)\left(\sum_{j=1}^{\infty}\sum_{t=M+1}^{N}p_{t}(N-t)(-1)^{j}\frac{t^{j}}{j!}\exp\left(-\frac{\pi tj}{\sqrt{6N}}\right)\right)\left(1+O\left(N^{-\frac{1}{2}}(\log N)^{4}\right)\right).$
Note
$\left|\sum_{j=1}^{\infty}\sum_{t=M+1}^{N}p_{t}(N-t)(-1)^{j}\frac{t^{j}}{j!}\exp\left(-\frac{\pi tj}{\sqrt{6N}}\right)\right|$
$\leq\left(\sum_{t=M+1}^{N}p_{t}(N-t)\right)\max_{\begin{subarray}{c}M+1\leq t\leq N\end{subarray}}\left|\sum_{j=1}^{\infty}(-1)^{j}\frac{t^{j}}{j!}\exp\left(-\frac{\pi tj}{\sqrt{6N}}\right)\right|$
$\leq p(N)\max_{\begin{subarray}{c}M+1\leq t\leq N\end{subarray}}\left(\exp\left(t\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)\right)-1\right)$
(4.7) $=O\left(\frac{p(N)}{(\log N)^{2}}\right).$
For the first part of (4), we use Lemma 3.8 at $t=T_{2}$ to obtain
(4.8) $p(N)-p_{T_{2}}(N)=\frac{2p(N)}{\log N}\left(1+O\left(\frac{1}{\sqrt{\log N}}\right)\right).$
The sum $\sum_{t=T_{2}+1}^{M}$ in (4) can be simplified using Abel summation
$\sum_{T_{2}<t\leq M}t^{j}\exp\left(-\frac{\pi t(j+1)}{\sqrt{6N}}\right)$
$=$ $M^{j}\frac{\exp\left(-\frac{\pi(j+1)}{\sqrt{6N}}\right)-\exp\left(-\frac{\pi(M+1)(j+1)}{\sqrt{6N}}\right)}{1-\exp\left(-\frac{\pi(j+1)}{\sqrt{6N}}\right)}$
@@
- $j\int_{T_{2}}^{M}u^{j-1}\frac{\exp\left(-\frac{\pi(j+1)}{\sqrt{6N}}\right)-\exp\left(-\frac{\pi(u+1)(j+1)}{\sqrt{6N}}\right)}{1-\exp\left(-\frac{\pi(j+1)}{\sqrt{6N}}\right)}du$
(4.9) $=$ $T_{2}^{j}\frac{\exp\left(-\frac{\pi(T_{2}+1)(j+1)}{\sqrt{6N}}\right)}{1-\exp\left(-\frac{\pi(j+1)}{\sqrt{6N}}\right)}-M^{j}\frac{\exp\left(-\frac{\pi(M+1)(j+1)}{\sqrt{6N}}\right)}{1-\exp\left(-\frac{\pi(j+1)}{\sqrt{6N}}\right)}$
$+$ $\frac{j}{1-\exp\left(-\frac{\pi(j+1)}{\sqrt{6N}}\right)}\int_{T_{2}}^{M}u^{j-1}\exp\left(-\frac{\pi(u+1)(j+1)}{\sqrt{6N}}\right)du.$
Pluging in $T_{2}$ from equation (4.2) in the first term of (4), we obtain
(4.10) $T_{2}^{j}\frac{\exp\left(-\frac{\pi(T_{2}+1)(j+1)}{\sqrt{6N}}\right)}{1-\exp\left(-\frac{\pi(j+1)}{\sqrt{6N}}\right)}=\frac{2}{(j+1)\log N}\left(1+O\left(\frac{\log\log N}{\log N}\right)\right).$
The second term of (4), evaluated at $M=N^{\frac{3}{4}-\delta}$, gives
(4.11) $M^{j}\frac{\exp\left(-\frac{\pi(M+1)(j+1)}{\sqrt{6N}}\right)}{1-\exp\left(-\frac{\pi(j+1)}{\sqrt{6N}}\right)}\leq\frac{C_{4}}{(j+1)N^{2}},$
for some positive constant $C_{4}$.
Combining (4.10), (4.11), we get the following expression for (4)
(4.12) $\sum_{T_{2}<t\leq M}t^{j}\exp\left(-\frac{\pi t(j+1)}{\sqrt{6N}}\right)=\frac{2}{(j+1)\log N}\left(1+O\left(\frac{\log\log N}{\log N}\right)\right)+E_{j},$
and again applying integration by parts we estimate $E_{j}$ as
$|E_{j}|\leq\frac{L}{(j+1)(\log N)^{2}},$
for some positive constant $L$ independent of $N$ and $j$.
Now the estimates from (4), (4.8) and (4.12) simplifies (4), and gives a lower bound for $A_{2}$
$A_{2}\geq$ $\left(\frac{2p(N)^{2}}{\log N}+\frac{2p(N)^{2}}{\log N}\sum_{j=1}^{\infty}\frac{(-1)^{j}}{(j+1)!}\right)\left(1+O\left(\frac{1}{\sqrt{\log N}}\right)\right)$
$=$ $\left(2-\frac{2}{e}\right)\frac{p(N)^{2}}{\log N}\left(1+O\left(\frac{1}{\sqrt{\log N}}\right)\right).$
Next, we calculate the lower bound for $A_{1}$.
Proof of Claim 2.
We again use (4.3) and (4.5), which are also valid for $T_{1}\leq t\leq T_{2}$, and obtain:
$A_{1}$ $=\sum_{t=T_{1}+1}^{T_{2}}c_{t}(N)p_{t}(N-t)$
$=\left(\sum_{t=T_{1}+1}^{T_{2}}c_{t}(N)p(N-t)\right)\left(1+O\left(\frac{1}{\sqrt{\log N}}\right)\right)$
(4.13) $\geq p(N)^{2}\left(\sum_{t=T_{1}+1}^{T_{2}}\exp\left(-t\exp\left(-\tfrac{\pi t}{\sqrt{6N}}\right)\right)\exp\left(-\frac{\pi t}{\sqrt{6N}}\right)\right)\left(1+O\left(\frac{1}{\sqrt{\log N}}\right)\right).$
@@
(4.14) $+G(T_{2})([T_{2}]-T_{2})-G(T_{1})([T_{1}]-T_{1}).$
After substituting the explicit values of $T_{1}$ and $T_{2}$ into the second term of the above expression, we get
$\int_{T_{1}}^{T_{2}}(t-[t])\,G^{\prime}(t)\,dt+G(T_{2})([T_{2}]-T_{2})-G(T_{1})([T_{1}]-T_{1})=O\!\left(N^{-\frac{1}{2}}\right).$
Choosing
$u(t):=\exp\left(-\frac{\pi t}{\sqrt{6N}}\right),$
@@
Since $F^{\prime}(u)>0$ for $u\in[u(T_{2}),u(T_{1})]$, the function $F(u)$ is increasing in this interval, and thus the integral dominates near the lower endpoint $u(T_{2})$. By the Laplace method for endpoint maxima [6], we obtain
@@
Now (4.15) gives an asymptotic formula for $\sum_{T_{1}<t\leq T_{2}}$ in (4). Substituting this into (4), we obtain
$A_{1}\geq\frac{2p(N)^{2}}{e\log N}\left(1+O\left(\frac{1}{\sqrt{\log N}}\right)\right).$
Finally, we obtain our required lower bound for $Z(N)$ from (4.1), using Claim 1 and Claim 2
$Z(N)\geq\frac{2\,p(N)^{2}}{\log N}\left(1+O\left(\frac{1}{\sqrt{\log N}}\right)\right).$
@@
By the Murnaghan-Nakayama rule (Theorem 3.1) and Lemma 3.2,
$Z_{t}(N)\geq c_{t}(N)\,p(N-t).$
From (1.1), the partition function p(Nt)p(N-t) is given by
$p(N-t)=\frac{1}{4(N-t)\sqrt{3}}\exp\left(\frac{2\pi}{\sqrt{6}}\sqrt{N-t}\right)\left(1+O\left((N-t)^{-1/2}\right)\right).$
Combining these bounds with Proposition 3.7, we complete the proof of Theorem 1.3. ∎